# Improved analysis of strong-interaction-stable doubly-bottom tetraquarks on the lattice

### v1

@@ I Introduction
The most deeply bound of the lattice results for the binding in the doubly bottom, $J^{P}=1^{+}$, $I=0$ and $I=1/2$ channels, those of Ref. Francis et al. (2017), came from an analysis using gauge-fixed wall sources and local sinks. Effective masses in such “wall-local” analyses typically approach the asymptotic ground state plateau from below, signalling negative excited state contamination. The fact that the ground state plateaus in this analysis were short, and reached at later Euclidean times, raised the possibility that the ground state masses were underestimated, and hence the associated bindings overestimated. This possibility was investigated in Ref. Hudspith et al. (2020), at a single $m_{\pi}$ ($\simeq 191$ MeV), using an improved “Box-Sink” construction designed to reduce excited-state contamination. The construction was shown to extend the ground state effective mass plateaus to much earlier Euclidean times and to significantly improve the ground state mass determinations. The results of that study confirmed that, at the single $m_{\pi}$ considered, the $I=0$ and $I=1/2$ binding energies were indeed overestimated in Ref. Francis et al. (2017). A similar conclusion was reached in Ref. Hudspith and Mohler (2023), also employing improved sink-side operators, designed to have better ground state overlap. For the $I=0$ channel, this conclusion is in keeping with the result obtained in Ref. Leskovec et al. (2019), which included two non-local “scattering operators” (products of two heavy-meson-like operators each projected to zero spatial momentum), though on the source side only, to avoid having to compute all-to-all propagators. These operators were expected to have improved overlap with the threshold $BB^{*}$ scattering state (presumably the first excited state in the channel), and hence, as a result, to also improve the ground-state signal. These expectations were born out by the pattern of relative overlaps of the source operators with the lowest few GEVP Eigenstates. Refs. Pflaumer et al. (2023); Wagner et al. (2023); Alexandrou et al. (2024) took further advantage of this improvement by including the non-local scattering operators at the sink as well as at the source, producing a reduced error on the ground-state $I=0$ binding energy. In this paper, we follow the alternate strategy of employing sink-side operators designed specifically to improved overlap with the ground state, extending the improved box-sink analysis of Ref. Hudspith et al. (2020) to ensembles with a range of $m_{\pi}$, allowing an extrapolation of the results for binding in the doubly bottom, $J^{P}=1^{+}$, $I=0$ and $I=1/2$ channels to physical $m_{\pi}$. Preliminary versions of this analysis were reported in Refs. Colquhoun et al. (2022a, b). We also revisit the variable heavy mass study of Ref. Francis et al. (2019) using the improved box-sink construction, and consider implications of the updated results for possible tetraquark binding in the $J^{P}=1^{+}$ bottom-charm channel.
@@ II Improvements to the previous analysis
Other improvements to our earlier analyses are as follows. In our initial study, Ref. Francis et al. (2017), we employed three $a\simeq 0.09$ fm, $32^{3}\times 64$ Wilson-clover ensembles, generated by the PACS-CS collaboration, having pion masses $\sim 164$, $299$ and $416$ MeV Aoki et al. (2009a, 2010). $m_{\pi}L$ is $>4$ for the latter two ensembles, but uncomfortably low ($\sim 2.4$) for the ensemble with the lightest pion mass. We also restricted ourselves to a two-operator basis, and hence a $2\times 2$ GEVP analyses. To ensure that potential finite-volume effects are under better control at lighter pion mass, and to improve the reliability of the extrapolation to physical $m_{\pi}$, we have generated additional, larger-volume, $48^{3}\times 64$ ensembles at the same lattice spacing, with pion masses $\sim 191$ and $\sim 165$ MeV corresponding to $m_{\pi}L=4.2$ and $m_{\pi}L=3.6$, respectively. To more fully map out the light-quark mass dependence of the doubly bottom channel binding energies, we have also carried out box-sink analyses of two additional PACS-CS ensembles with heavier pion masses, $\sim 575$ and $707$ MeV, in addition to the box-sink updates of our earlier analyses of the three previously studied $32^{3}\times 64$ PACS-CS ensembles. Finally, to help better determine the ground-state signal, we have expanded our operator bases beyond the two-operator basis used in Francis et al. (2017), and hence the number of correlators entering our analyses, and we employ the box-sink analysis described above (Eq. (2)) to all operators. We also perform multi-exponential fits in place of the GEVP previously used, which will be explained in detail below.
Table 1 provides details of the ensembles employed in this study, while further details of the operator bases used in the various channels are discussed below. Pion masses in Table 1 are obtained from multi-exponential fits, agreeing well with previous values obtained on these ensembles Aoki et al. (2009b).
@@ III.1 Ensembles and correlator calculation
This same set (with $b$ replaced by $h$) is used for the local operators of the double-variable-heavy-mass $T_{\operatorname{ud\bar{h}\bar{h}}}$ and $T_{\operatorname{ls\bar{h}\bar{h}}}$ channels. The operators for the $I(J^{P})=0(1^{+})/0(0^{+})$ and $\frac{1}{2}(1^{+})/\frac{1}{2}(0^{+})$ channels having one physical-b-mass and one variable-heavy-mass antiquark, $T_{\operatorname{ud\bar{b}\bar{h}}}$ and $T_{\operatorname{ls\bar{b}\bar{h}}}$, are obtained from the “$udcb$” and “$uscb$” rows of the same table by replacing the “$c$” (which in Ref. Hudspith et al. (2020) denoted a relativistic charm antiquark) with $h$ (denoting here a NRQCD heavy antiquark whose mass we allow to vary over a range including values both heavier and lighter than physical $m_{b}$). [^2] With NRQCD used for the heavy quarks, the variable heavy mass, $m_{h}$, must, of course, remain in the region of validity of the use of NRQCD. The range thus does not extend down to $m_{h}$ as low as physical $m_{c}$.
For all ensembles except E7 we have, for each sink operator, four box radii squared [^3] Owing to limited resources, for E7 we only have two box radii squared.. In lattice units, these are in the range 30-49 for the tetraquarks, and 16-49 for the mesons (for which we will only use one source and sink operator per box radius). Our strategy is to use multiple sinks in a simultaneous fit to reliably extract the ground- and excited-state energies, as will be detailed below.
@@ V Summary and discussion
In this paper we have revisited the earlier lattice determination of the binding energies of the flavour-antitriplet of doubly-bottom $J^{P}=1^{+}$ tetraquarks reported in Ref. Francis et al. (2017), introducing a number of significant improvements and producing results which update and supersede those of Ref. Francis et al. (2017). Relative to the previous study, the new analysis (i) employs expanded sets of interpolating operators for the $\operatorname{ud\bar{b}\bar{b}}$ and $\operatorname{ls\bar{b}\bar{b}}$ channels, (ii) adds two newly generated, larger-volume ensembles with $m_{\pi}<200$ MeV, aimed at reducing possible FV effects, and (iii) rather than using local operators at the sink, as done in Ref. Francis et al. (2017), implements the improved box-sink construction of Ref. Hudspith et al. (2020). As shown in Ref. Hudspith et al. (2020), this construction significantly reduces excited state contamination of the ground-state signal, leading to much longer ground-state effective-mass plateaus, beginning at much earlier Euclidean times, than those obtained in the local-sink analysis of Ref. Francis et al. (2017).
@@
Our result for the $I=0$, $J^{P}=1^{+}$ $\operatorname{ud\bar{b}\bar{b}}$ binding, $115(17)$ MeV, shows some tension with the results of two earlier studies, Refs. Francis et al. (2017) and  Mohanta and Basak (2020), which found $189(10)(3)$ MeV and $187(18)$ MeV, respectively, though, as noted above, the former result is superseded by that of the current, improved analysis. Our new result agrees within errors with all other recent lattice results for this channel, with Ref. Junnarkar et al. (2019) finding $143(34)$ MeV, Ref. Leskovec et al. (2019) $128(24)(10)$ MeV, Ref. Hudspith and Mohler (2023) $112(13)$ MeV, Ref. Aoki et al. (2023) $83(10)(20)$ MeV and Ref. Alexandrou et al. (2024) $100(10)\left({}^{+36}_{-43}\right)$ MeV. It is worth noting that among these the physical-point HALQCD result, $83(10)(20)$ MeV Aoki et al. (2023), is obtained from a significantly longer physical-point extrapolation (of results with $m_{\pi}\sim 410$, $575$ and $700$ MeV) than are the results of Refs. Junnarkar et al. (2019); Leskovec et al. (2019); Hudspith and Mohler (2023); Alexandrou et al. (2024) or the current analysis, all of which include ensembles with $m_{\pi}$ at least as low as $\sim 220$ MeV. A potentially large systematic error can arise from the use of NRQCD. It cannot be estimated within our work since we have just one lattice spacing, but a previous study Hudspith and Mohler (2022) has found it to be of the order of 10 MeV.

### v2

@@ I Introduction
The most deeply bound of the lattice results for the binding in the doubly bottom, $J^{P}=1^{+}$, $I=0$ and $I=1/2$ channels, those of Ref. Francis et al. (2017), came from an analysis using gauge-fixed wall sources and local sinks. Effective masses in such “wall-local” analyses typically approach the asymptotic ground state plateau from below, signalling negative excited state contamination. The fact that the ground state plateaus in this analysis were short, and reached at later Euclidean times, raised the possibility that the ground state masses were underestimated, and hence the associated bindings overestimated. This possibility was investigated in Ref. Hudspith et al. (2020), at a single $m_{\pi}$ ($\simeq 191$ MeV), using an improved “Box-Sink” construction designed to reduce excited-state contamination. The construction was shown to extend the ground state effective mass plateaus to much earlier Euclidean times and to significantly improve the ground state mass determinations. The results of that study confirmed that, at the single $m_{\pi}$ considered, the $I=0$ and $I=1/2$ binding energies were indeed overestimated in Ref. Francis et al. (2017). A similar conclusion was reached in Ref. Hudspith and Mohler (2023), also employing improved sink-side operators, designed to have better ground state overlap. For the $I=0$ channel, this conclusion is in keeping with the result obtained in Ref. Leskovec et al. (2019), which included two non-local “scattering operators” (products of two heavy-meson-like operators each projected to zero spatial momentum), though on the source side only, to avoid having to compute all-to-all propagators. These operators were expected to have improved overlap with the threshold $BB^{*}$ scattering state (presumably the first excited state in the channel), and hence, as a result, to also improve the ground-state signal. These expectations were born out by the pattern of relative overlaps of the source operators with the lowest few GEVP Eigenstates. Refs. Pflaumer et al. (2023); Wagner et al. (2023); Alexandrou et al. (2024) took further advantage of this improvement by including the non-local scattering operators at the sink as well as at the source, producing a reduced error on the ground-state $I=0$ binding energy. In this paper, we follow the alternate strategy of employing sink-side operators designed specifically to improve overlap with the ground state, extending the improved box-sink analysis of Ref. Hudspith et al. (2020) to ensembles with a range of $m_{\pi}$, allowing an extrapolation of the results for binding in the doubly bottom, $J^{P}=1^{+}$, $I=0$ and $I=1/2$ channels to physical $m_{\pi}$. Preliminary versions of this analysis were reported in Refs. Colquhoun et al. (2022a, b). We also revisit the variable heavy mass study of Ref. Francis et al. (2019) using the improved box-sink construction, and consider implications of the updated results for possible tetraquark binding in the $J^{P}=1^{+}$ bottom-charm channel.
@@ II.1 The box sink construction
II.2 Increased number of ensembles
Another improvement to our earlier analysis is the inclusion of a larger number of ensembles, with different pion masses, including two newly generated ensembles.
In our initial study, Ref. Francis et al. (2017), we employed three $a\simeq 0.09$ fm, $32^{3}\times 64$ Wilson-clover ensembles, generated by the PACS-CS collaboration, having pion masses $\sim 164$, $299$ and $416$ MeV Aoki et al. (2009a, 2010). $m_{\pi}L$ is $>4$ for the latter two ensembles, but uncomfortably low ($\sim 2.4$) for the ensemble with the lightest pion mass. To ensure that potential finite-volume effects are under better control at lighter pion mass, and to improve the reliability of the extrapolation to physical $m_{\pi}$, we have generated additional, larger-volume, $48^{3}\times 64$ ensembles at the same lattice spacing, with pion masses $\sim 191$ and $\sim 165$ MeV corresponding to $m_{\pi}L=4.2$ and $m_{\pi}L=3.6$, respectively. To more fully map out the light-quark mass dependence of the doubly bottom channel binding energies, we have also carried out box-sink analyses of two additional PACS-CS ensembles with heavier pion masses, $\sim 575$ and $707$ MeV, in addition to the box-sink updates of our earlier analyses of the three previously studied $32^{3}\times 64$ PACS-CS ensembles. Details of all ensembles used are given in the next Section III.1, and they are listed in Table 1.
II.3 Operator basis and multi-exponential fits
Finally, to help better determine the ground-state signal, we have expanded our operator bases beyond the two-operator basis used in Francis et al. (2017), and hence the number of correlators entering our analyses. We also apply the box-sink analysis described above (Eq. (2)) to all operators, leading to an increased operator basis, as for each operator, we simultaneously fit data with several box sink radii. Further details of the operator bases used in the various channels are discussed in Appendix A.
We also perform multi-exponential fits in place of the GEVP previously used, which will be explained in detail in Section III.2. This is not necessarily an improvement in itself, but rather a different approach to the problem, as the GEVP and multi-exponential fits are both equally valid methods to extract energies from correlator data. One potential advantage of the multi-exponential fit is that it does not require a complete matrix of all source sink combinations, however, in our case we include all combinations anyway. Another is that for a tower of very finely spaced excited states, the fit can more easily approximate the areas of spectral density over a number of states. Pion masses in Table 1 are also obtained from multi-exponential fits, agreeing well with previous values obtained on these ensembles Aoki et al. (2009b).
@@ III.1 Ensembles and correlator calculation
This same set (with $b$ replaced by $h$) is used for the local operators of the double-variable-heavy-mass $T_{\operatorname{ud\bar{h}\bar{h}}}$ and $T_{\operatorname{ls\bar{h}\bar{h}}}$ channels. The operators for the $I(J^{P})=0(1^{+})/0(0^{+})$ and $\frac{1}{2}(1^{+})/\frac{1}{2}(0^{+})$ channels having one physical-b-mass and one variable-heavy-mass antiquark, $T_{\operatorname{ud\bar{b}\bar{h}}}$ and $T_{\operatorname{ls\bar{b}\bar{h}}}$, are obtained from the “$udcb$” and “$uscb$” rows of the same table by replacing the “$c$” (which in Ref. Hudspith et al. (2020) denoted a relativistic charm antiquark) with $h$ (denoting here a NRQCD heavy antiquark whose mass we allow to vary over a range including values both heavier and lighter than physical $m_{b}$). [^2] With NRQCD used for the heavy quarks, the variable heavy mass, $m_{h}$, must, of course, remain in the region of validity of the use of NRQCD. The range thus does not extend down to $m_{h}$ as low as physical $m_{c}$. Appendix A lists the operators used in our case explicitly. For all ensembles except E7 we have, for each sink operator, four box radii squared [^3] Owing to limited resources, for E7 we only have two box radii squared.. In lattice units, these are in the range 30-49 for the tetraquarks, and 16-49 for the mesons (for which we will only use one source and sink operator per box radius). Our strategy is to use multiple sinks in a simultaneous fit to reliably extract the ground- and excited-state energies, as will be detailed below.
@@ V Summary and discussion
In this paper we have revisited the earlier lattice determination of the binding energies of the flavour-antitriplet of doubly-bottom $J^{P}=1^{+}$ tetraquarks reported in Ref. Francis et al. (2017), introducing a number of significant improvements and producing results which update and supersede those of Ref. Francis et al. (2017). Relative to the previous studies, our new analysis employs several improvements. First, rather than using local operators at the sink, as done in Ref. Francis et al. (2017), we implement the improved box-sink construction of Ref. Hudspith et al. (2020). As shown in Ref. Hudspith et al. (2020), this construction significantly reduces excited state contamination of the ground-state signal, leading to much longer ground-state effective-mass plateaus, beginning at much earlier Euclidean times, than those obtained in the local-sink analysis of Ref. Francis et al. (2017). Second, we add two newly generated, larger-volume ensembles with $m_{\pi}<200$ MeV, aimed at reducing possible FV effects. Finally, we employ an expanded sets of interpolating operators for the $\operatorname{ud\bar{b}\bar{b}}$ and $\operatorname{ls\bar{b}\bar{b}}$ channels, and approach the correlator fits using a multi-exponential fit, rather than a GEVP.
It is difficult to separate the effects of these different changes out fully, but it is clear from Hudspith et al. (2020) that the box sink method is likely the main driver for the shallower bindings we find with respect to previous analyses using local sinks. Whilst the additional pion masses certainly improved our uncertainties and confidence in our findings, it is clear from Fig. 4 that without these additional data points our results would be broadly similar. Similarly, for the expanded operator basis and multi-exponential fits, it is more the case that these changes give us enhanced confidence in our results, than are directly responsible for any explicit shift in values from earlier analyses.
@@
Our result for the $I=0$, $J^{P}=1^{+}$ $\operatorname{ud\bar{b}\bar{b}}$ binding, $115(17)$ MeV, shows some tension with the results of two earlier studies, Refs. Francis et al. (2017) and  Mohanta and Basak (2020), which found $189(10)(3)$ MeV and $187(18)$ MeV, respectively, though, as noted above, the former result is superseded by that of the current, improved analysis. Our new result agrees within errors with all other recent lattice results for this channel, with Ref. Junnarkar et al. (2019) finding $143(34)$ MeV, Ref. Leskovec et al. (2019) $128(24)(10)$ MeV, Ref. Hudspith and Mohler (2023) $112(13)$ MeV, Ref. Aoki et al. (2023) $83(10)(20)$ MeV and Ref. Alexandrou et al. (2024) $100(10)\left({}^{+51}_{-36}\right)$ MeV. It is worth noting that among these the physical-point HALQCD result, $83(10)(20)$ MeV Aoki et al. (2023), is obtained from a significantly longer physical-point extrapolation (of results with $m_{\pi}\sim 410$, $575$ and $700$ MeV) than are the results of Refs. Junnarkar et al. (2019); Leskovec et al. (2019); Hudspith and Mohler (2023); Alexandrou et al. (2024) or the current analysis, all of which include ensembles with $m_{\pi}$ at least as low as $\sim 220$ MeV. A potentially large systematic error can arise from the use of NRQCD. It cannot be estimated within our work since we have just one lattice spacing, but a previous study Hudspith and Mohler (2022) has found it to be of the order of 10 MeV.
@@ Acknowledgements
Appendix A Operator basis
The operator basis used in this work is a subset of that used in Hudspith et al. (2020). Copying the generalised notation in that reference for our specific case, our operators are,
$\begin{gathered}D(\Gamma_{1},\Gamma_{2})=(\psi_{a}^{T}C\Gamma_{1}\phi_{b})(\bar{\theta}_{a}C\Gamma_{2}\bar{\omega}^{T}_{b}),\\ E(\Gamma_{1},\Gamma_{2})=(\psi_{a}^{T}C\Gamma_{1}\phi_{b})(\bar{\theta}_{a}C\Gamma_{2}\bar{\omega}^{T}_{b}-\bar{\theta}_{b}C\Gamma_{2}\bar{\omega}^{T}_{a}),\\ M(\Gamma_{1},\Gamma_{2})=(\bar{\theta}\Gamma_{1}\psi)(\bar{\omega}\Gamma_{2}\phi),\qquad N(\Gamma_{1},\Gamma_{2})=(\bar{\theta}\Gamma_{1}\phi)(\bar{\omega}\Gamma_{2}\psi).\end{gathered}$ (12)
The list of operators used in this work for each tetraquark channel is given in Tables 4 and 5.
Type $(\psi\phi\theta\omega)$ $I(J)^{P}$ Diquark-Antidiquark Dimeson
$ud\bar{b}\bar{h}$ $0(1)^{+}$ $E(\gamma_{5},\gamma_{i}),E(\gamma_{t}\gamma_{5},\gamma_{i}\gamma_{t})$ $M(\gamma_{5},\gamma_{i})-N(\gamma_{5},\gamma_{i})$
$M(I,\gamma_{i}\gamma_{5})-N(I,\gamma_{i}\gamma_{5})$
$ud\bar{h}\bar{h}$ $0(1)^{+}$ $D(\gamma_{5},\gamma_{i}),\;D(\gamma_{t}\gamma_{5},\gamma_{i}\gamma_{t})$ $M(\gamma_{5},\gamma_{i})-N(\gamma_{5},\gamma_{i})$
$M(I,\gamma_{i}\gamma_{5})-N(I,\gamma_{i}\gamma_{5})$
$ls\bar{h}\bar{h}$ $\frac{1}{2}(1)^{+}$ $D(\gamma_{5},\gamma_{i}),\;D(\gamma_{t}\gamma_{5},\gamma_{i}\gamma_{t})$ $M(\gamma_{5},\gamma_{i}),\;M(I,\gamma_{i}\gamma_{5})$
$N(\gamma_{5},\gamma_{i}),\;N(I,\gamma_{i}\gamma_{5})$
$ls\bar{b}\bar{h}$ $\frac{1}{2}(1)^{+}$ $E(\gamma_{5},\gamma_{i}),\;E(\gamma_{t}\gamma_{5},\gamma_{i}\gamma_{t})$ $M(\gamma_{5},\gamma_{i}),\;M(I,\gamma_{i}\gamma_{5})$
$N(\gamma_{5},\gamma_{i}),\;N(I,\gamma_{i}\gamma_{5})$
Table 4: $J^{P}=1^{+}$ operators (from Eq. (12)) used in this work.
Type $(\psi\phi\theta\omega)$ $I(J)^{P}$ Diquark-Antidiquark Dimeson
$ud\bar{b}\bar{h}$ 0($0(0)^{+}$ $E(\gamma_{5},\gamma_{5}),E(\gamma_{t}\gamma_{5},\gamma_{t}\gamma_{5})$ $M(\gamma_{5},\gamma_{5})-N(\gamma_{5},\gamma_{5})$
$M(I,I)-N(I,I)$
$ls\bar{b}\bar{h}$ $\frac{1}{2}(0)^{+}$ $E(\gamma_{5},\gamma_{5}),\;E(\gamma_{t}\gamma_{5},\gamma_{t}\gamma_{5})$ $M(\gamma_{5},\gamma_{5}),\;M(I,I)$
$N(\gamma_{5},\gamma_{5}),\;N(I,I)$
Table 5: $J^{P}=0^{+}$ operators (from Eq. (12)) used in this work.